# Shaocong Xu1,2, Pengfei Li2, Xinyu Liu2, Qianpu Sun2, Yang Li2, Shihui Guo1, Zhen Wang3, Bo Jiang3, Rui Wang3, Kehua Sheng3, Bo Zhang3, Hao Zhao2

### v1

@@ Abstract
LiDAR-based semantic scene understanding is an important module in the modern autonomous driving perception stack. However, identifying Out-Of-Distribution (OOD) points in a LiDAR point cloud is challenging as point clouds lack semantically rich features when compared with RGB images. We revisit this problem from the perspective of selective classification, which introduces a selective function into the standard closed-set classification setup. Our solution is built upon the basic idea of abstaining from choosing any known categories but learns a point-wise abstaining penalty with a margin-based loss. Synthesizing outliers to approximate unlimited OOD samples is also critical to this idea, so we propose a strong synthesis pipeline that generates outliers originated from various factors: unrealistic object categories, sampling patterns and sizes. We demonstrate that learning different abstaining penalties, apart from point-wise penalty, for different types of (synthesized) outliers can further improve the performance. We benchmark our method on SemanticKITTI and nuScenes and achieve state-of-the-art results. Risk-coverage analysis further reveals intrinsic properties of different methods. Codes and models will be publicly available.
@@ Introduction
3D LiDAR sensors have been widely adopted in modern autonomous cars, primarily due to their capability to directly capture precise and long-range distance measurements of the surrounding environment, when compared with RGB cameras. However, LiDARs are not perfect and a notable disadvantage is the lack of semantically rich features. Human eyes can easily recognize semantic categories from an RGB image while the task becomes substantially more difficult when the input is a LiDAR point cloud. This conclusion also holds for machine eyes (i.e., computer vision algorithms), as evidenced by the fact that LiDAR-based semantic segmentation usually under-performs its RGB-based counterpart.
The lack of semantically rich features also negatively impacts LiDAR-based road anomaly detection. Road anomaly detection is an extension of semantic segmentation, which aims to identify regions not belonging to the pre-defined label set. RGB-based road anomaly detection draws a lot of attention recently from both academia and industry and sees remarkable progress (Tian et al. 2022; Liu et al. 2022; Tian et al. 2023). By contrast, LiDAR-based anomaly detection (Cen et al. 2022) is still in a very early stage and reports poor quantitative results on public benchmarks. The focus of this study is to address this important yet challenging problem.
The aforementioned difficulty is demonstrated through Fig. 1-(a). As shown in the upper panel, although this is an extremely difficult case in which the anomaly is very small, we can still find the plank left on the road. Meanwhile, the SOTA algorithm successfully highlights it in the heatmap. On the contrary, detecting anomaly in a point cloud is much more challenging due to the lack of semantically rich features, shown in the lower panel of Fig. 1-(a). There is some furniture on the road, which is clearly anomalous, but both human eyes and computer vision algorithms struggle to identify them.
Figure 1: a. 3D anomaly detection is much more difficult than its 2D counterpart; b. Comparison of our ShapeNet-based synthesis pipeline and the former resizing-based synthesis pipeline; c. Comparison of our method with REAL in SemanticKITTI using Risk-Coverage, AUPR-Coverage, and AUROC-Coverage curves. Our method is better.
How to capture the subtle differences between known and anomalous classes in a LiDAR point cloud? We make contributions from both the perspective of formulation and data.
Formulation. REAL (Cen et al. 2022) exploits dummy classifiers to aggregate information in anomalous objects, which is spiritually similar to Selective Classification (SC) (Feng et al. 2023). We reformulate the point cloud anomaly detection problem strictly using the principle of SC. What’s more, we learn the abstaining penalty in a point-wise manner, which can better capture the subtle difference in point clouds. Meanwhile, our formulation also learns different penalties for different data sources. Finally, we go back to the origin of SC and delve into the risk-coverage analysis. As illustrated in Fig. 1-(c), our method exhibits a significantly better performance compared to REAL across various levels of coverage, demonstrating its superiority. Through the lens of risk-coverage curve, we attain a deeper understanding of the performance gains achieved by our approach, and insights into the nature of different methods.
Data. The SOTA method (Cen et al. 2022) synthesizes pseudo anomalous objects by resizing known objects, as shown in the upper panel of Fig. 1-(b). However, we observe that pseudo anomalous objects synthesized in this way fail to well represent the long-tail distribution of road anomalies, in two regards. Firstly, objects from existing road scene understanding datasets are limited in category and thus resizing them hardly compensates for the lack of semantically rich features. Secondly, naive resizing violates the sampling pattern of real LiDAR sensors: points on enlarged objects get sparser while those on shrunk objects get denser. This leads to a shortcut problem: the model may find a trivial solution to distinguish pseudo anomalous objects from known objects solely using sparsity.
To generate better pseudo anomalous objects, we propose to introduce objects from an external dataset, ShapeNet (Chang et al. 2015), into existing scenes. Specifically, we merge randomly selected ShapeNet objects into road scenes by changing the radii of certain scene points in spherical coordinates. By replacing the radii, we preserve the sampling pattern of realistic point clouds as shown in the lower panel of Fig. 1-(b). We compensate for the lack of semantically rich features in point clouds by incorporating semantically rich anomalous objects, as ShapeNet covers a wide spectrum of categories.
@@
  • We propose a new approach for endowing 3D segmentation models with the capability of detecting anomalies, which can be utilized in conjunction with existing segmentation models.
  • We utilize ShapeNet object as a guide to modify the radii of existing points to synthesize more realistic and more diverse anomalies.
  • We incorporate risk-coverage curve metrics into our evaluation process. By utilizing risk-coverage curve analysis, we are able to identify the nature of improvements and gain valuable insights into the strengths and weaknesses of different approaches.
@@
  • Our proposed method has achieved new SOTA performance for anomaly detection not only in previously established anomaly detection metrics, but also in the newly introduced risk-coverage curve metric, on SemanticKITTI and NuScenes.
Selective Classification Problem Formulation
@@
The seminal point cloud anomaly detection method REAL (Cen et al. 2022) lacks a principled framework, hindering both theoretical and experimental analysis. So we first formalize the definition of SC and put different methods under a unified theoretical lens. Experimentally, the risk-coverage trade-off analysis that comes along with this framework, allows us to reveal in-depth differences between methods.
@@
In SC, our goal is to learn predictive models that know what they do not know or when they should abstain [^1] The abstaining penalty will be defined later. from making decisions. Here we consider a generic SC definition, which is agnostic of network and application. A supervised classification task is formulated as follows. Let $\mathcal{X}$ be any feature space and $\mathcal{Y}$ a label space. In point cloud anomaly detection, $\mathcal{X}$ could be point clouds, and $\mathcal{Y}$ could be class labels [^2] These class labels include an outlier label. of each point cloud. Let $P(\mathcal{X},\mathcal{Y})$ be a distribution over $\mathcal{X}\times\mathcal{Y}$. A model $f:\mathcal{X}\rightarrow\mathcal{Y}$ is called a prediction function and its true risk used to evaluate the performance of $f$ w.r.t. $P$ is $R(f):=E_{P(\mathcal{X},\mathcal{Y})}[\ell(f(\mathbf{x}),\mathbf{y})]$, where $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a given loss function, for example, the Cross Entropy (CE) loss. Given a labeled set $S_{m}=\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{m}\subseteq(\mathcal{X}\times\mathcal{Y})^{m}$ sampled i.i.d. from $P(X,Y)$, the empirical risk of the classifier $f$ is $\hat{r}(f|S_{m}):=\frac{1}{m}\sum_{i=1}^{m}\ell(f(\mathbf{x_{i}}),\mathbf{y_{i}})$.
@@
Remark. Later we will put different methods under the unified lens of SC and use the risk-coverage curve to probe their intrinsic properties.
Finally, we clarify that the continuous risk and coverage defined above are empirically calculated using a fixed set in practice. For any given labeled set $S_{m}$, the empirical selective risk is
@@
Related Works
Anomaly Detection in Autonomous Driving. Anomaly detection in autonomous driving is vital for ensuring the safety of ego-cars in open-world environments by identifying OOD objects. Extensive research has been conducted in 2D perception, specifically with semantic segmentation models. Unsupervised methods (Jung et al. 2021; Bevandić et al. 2021) involve post-processing predicted logits from frozen segmentation models to detect anomalies. Supervised methods (Tian et al. 2022; Grcić, Bevandić, and Šegvić 2022; Chan, Rottmann, and Gottschalk 2021) utilize auxiliary datasets like COCO (Lin et al. 2014) to synthesize anomaly objects in training images (e.g., Cityscapes (Cordts et al. 2016)) and retrain the segmentation model using outlier exposure (Hendrycks, Mazeika, and Dietterich 2018).
While significant advancements have been made in 2D anomaly detection, the exploration in the context of 3D point clouds remains limited. Cen et al. (2022) use resizing-based synthesis pipeline and calibration loss to achieve the discrimination of anomaly points. However, their focus primarily revolves around the open-world segmentation setting without extensively analyzing calibration effectiveness. Li and Dong (2023) propose an adversarial prototype framework that improves performance but involves complex network design and computationally expensive training. Considering these limitations, we present our method and substantiate its effectiveness through extensive experiments.
Selective Classification. SC can be broadly classified into two groups: 1) the first group focuses on addressing SC through the use of additional heads/logits (Geifman and El-Yaniv 2017, 2019; Liu et al. 2019; Feng et al. 2023; Gal and Ghahramani 2016; Chow 1970); 2) the second group tackles SC through cost-sensitive classification techniques (Charoenphakdee et al. 2021; Mozannar and Sontag 2020). Motivated by the extra head/logits design, resembling anomaly detection, we design new anomaly detector from the perspective of SC and utilize SC’s evaluation metrics as alternative performance measures for our anomaly detector.
Synthetic Data. Due to the limited availability and high cost associated with acquiring high-quality real-world datasets, synthetic data has been extensively employed in the field of scene understanding (Su et al. 2015; Movshovitz-Attias, Kanade, and Sheikh 2016; Zhang et al. 2017b; Handa et al. 2016; McCormac et al. 2017; Zhang et al. 2017a; Song et al. 2017). In this work, we introduce ShapeNet objects into the existing scene to effectively compensate for the lack of semantically rich features in point clouds.
@@
Figure 2: Our method’s pipeline: LiDAR point cloud input is processed by a feature extractor to extract features. These features are then used by the inlier and outlier classifiers to predict class logits. Point-wise abstain and penalty loss are calculated based on these predictions.
3D semantic segmentation is the task of assigning a class label to each point in a given LiDAR point cloud. Anomaly detection, on the other hand, is an extension to semantic segmentation, which aims to identify points that do not belong to the predefined inlier label set (learning to detect anomalous outliers). While the seminal work (Cen et al. 2022) proposes a viable solution REAL, we revisit it using the unified theoretical framework of SC and propose ours that effectively addresses the lack of semantically rich features.
Apart from the formulation contribution, we also make a contribution from the perspective of data. We design a novel anomaly synthesis pipeline that leverages the richness of the ShapeNet and adheres to the realistic LiDAR distribution to compensate for the lack of semantically rich features.
Architecture Overview
As shown in Fig. 2, the input LiDAR point cloud $\mathbf{x}\in\mathbb{R}^{n\times 3}$, sampled from $S_{m}$, is denoted on the left with ShapeNet objects integrated into it. Then, $\mathbf{x}$ is fed into the feature extractor $\psi$ followed by a inlier classifier $f$ to predict the inlier logit $\mathbf{\hat{y}}\in\mathbb{R}^{n\times c}$, where $c$ is the number of inlier classes. Additionally, an outlier classifier $g$ is employed to predict the outlier logit $\mathbf{\hat{o}}\in\mathbb{R}^{n\times 1}$. As such, $(f,g)$ instantiates a selective model mentioned above. These operations can be expressed as follows:
$\begin{gathered}\mathbf{\hat{y}}:=f(\psi(\mathbf{x}))\\ \mathbf{\hat{o}}:=g(\psi(\mathbf{x}))\\ \mathbf{\tilde{y}}:=[\mathbf{\hat{y}},\mathbf{\hat{o}}]:=\Biggl{\{}\mathbf{\tilde{y}_{i}}:=[\hat{y}_{i},\hat{o}_{i}]\Big{|}i=1,\ldots,n\Biggr{\}}\\ \mathbf{p}:=\Biggl{\{}p_{i,j}=\frac{e^{\tilde{y}_{i,j}}}{\sum_{k=1}^{c+1}e^{\tilde{y}_{i,k}}}\Big{|}i=1,\ldots,n;j=1,\ldots,c+1\Biggr{\}}\\ \mathbf{p^{y}},\mathbf{p^{o}}:=\mathbf{p}\end{gathered}$ (6)
@@
Cen et al. (2022) introduce the first approach REAL to conduct anomaly detection in LiDAR point clouds. Their added dummy classifiers can also be thought of as $g$, under the SC framework, but there is a key difference. They observe that numerous real outliers are wrongly classified as inlier classes with high probabilities. To address this issue, they propose a Calibration Cross Entropy (CCE) loss function to drive the outlier logit of the inlier sample to the second largest. We formalize [^3] This is equivalent to the loss used in the REAL paper although REAL describes it in a different manner. this loss as follows:
$\begin{gathered}\ell:=\frac{1}{m}\sum_{S_{m}}\frac{1}{n}\sum_{i=1}^{n}\Biggl{\{}\underbrace{-\log{\frac{e^{\tilde{y}_{i,y_{i}}}}{\sum_{k=1}^{c+1}e^{\tilde{y}_{i,k}}}}}_{\text{CE}}\\ \underbrace{-\lambda\mathbb{I}(y_{i}\neq c+1)\log{\frac{e^{\tilde{y}_{i,c+1}}}{\sum_{k=1\&k\neq y_{i}}^{c+1}e^{\tilde{y}_{i,k}}}}}_{\text{CCE}}\Biggr{\}}\\ \end{gathered}$ (7)
Here, $y_{i}\in\{1,\ldots,c,c+1\}$ signifies the ground truth corresponding to $x_{i}\in\mathbf{x}$, where $\{1,\ldots,c\}$ represents inlier class labels while $\{c+1\}$ indicates the outlier class label. $\mathbf{y}$ and $\mathbf{x}$ are sampled i.i.d. from $S_{m}$. $\mathbb{I}(\cdot)$ is the indicator function. $\lambda$ is a hyperparameter. A notable fact is that REAL does not provide an ablation study for this CCE loss and as shown in  Tab. 1, removing this CCE loss can indeed significantly improve standard anomaly detection metrics. But why this happens cannot be understood through metrics like AUPR or AUROC, highlighting the need to revisit REAL under the lens of SC. Specifically, under the SC perspective, we can use the risk-coverage analysis to investigate it.
Table 1: Removing CCE in REAL can actually improve the anomaly detection metrics on SemanticKITTI.
CE CCE AUPR AUROC $\text{mIoU}_{\text{old}}$
20.00 84.90 57.80
26.68 87.60 58.28
A Closer Look
@@
Figure 3: Statistics of outlier probability $\mathbf{p^{o}}$ for inlier and outlier samples under different settings of REAL on SemanticKITTI. For inlier samples, we woule like to observe more samples with $\mathbf{p^{o}}\in[0,0.1]$, while for outlier samples, less samples with $\mathbf{p^{o}}\in[0,0.1]$ is desirable.
@@
Figure 4: Anomaly synthesis pipeline.
Point-wise Abstaining Penalty Training Paradigm
Now that we have taken the perspective of SC, we argue that making the selective function $g$ better is the key factor. Recall that our architecture (Fig. 2) directly predicts the outlier probability $\mathbf{p^{o}}$, on which we can impose different thresholds to achieve different coverage. Our proposal is to replace REAL’s problematic loss with a new one that imposes abstaining penalty on the outlier probability $\mathbf{p^{o}}$:
@@
The $\alpha$ value denotes the point-wise penalty that controls the tendency to abstain or not. Since this value is defined in a point-wise manner, it has the potential to capture the subtle difference between inliers and outliers despite point clouds lack semantically rich features for anomaly detection. Collaborating with the above point-wise abstain loss (right-top in Fig. 2), we introduce a point-wise penalty loss (right-bottom in Fig. 2), which is expressed as follows:
$\begin{gathered}\mathbf{\alpha}:=\left\{\alpha_{i}=-\log(\sum_{j=1}^{c}e^{\hat{y}_{i,j}})|i=1,\ldots,n\right\}\\ \ell^{\text{penalty}}:=\frac{1}{m}\sum_{S_{m}}\frac{1}{n}\sum_{i=1}^{n}\Biggl{\{}\mathbb{I}(y_{i}\neq c+1)\text{max}(\alpha_{i}-m_{\text{in}},0)\\ +\mathbb{I}(y_{i}=c+1)\text{max}(m_{\text{out}}-\alpha_{i},0)\Biggr{\}}\end{gathered}\vspace{-0.5em}$ (9)
Figure 5: Left: Relationship between point-wise penalty loss and penalty $\alpha$. Right: Comparison with different anomaly synthesis pipeline using the risk-coverage curve.
As shown in Fig. 5-left, the hyperparameters $m_{\text{in}}$ and $m_{\text{out}}$ ensure that the inlier samples are associated with penalties lower than $m_{\text{in}}$, while the outlier samples exhibit penalties higher than $m_{\text{out}}$. Note that in our experiments, the penalties are negative as we set the value of $m_{\text{in}}$ to -12, and $m_{\text{out}}$ to -6. Intuitively, minimizing the point-wise penalty loss assigns lower $\alpha_{i}$ penalties to inlier samples. This results in higher $(-\alpha_{i})^{2}$, requiring a higher $p^{y}_{i,y_{i}}$ to achieve lower point-wise abstain loss, which is desirable for inliers. Likewise, higher $\alpha_{i}$ penalties are allocated to outlier samples to minimize point-wise penalty loss. This leads to lower $(-\alpha_{i})^{2}$, requiring a lower $\{p^{y}_{i,j}|j=1,\ldots,c\}$ to achieve lower point-wise abstain loss.
As such, the total loss can be expressed as follows:

### v3

# LiON: Learning Point-wise Abstaining Penalty for LiDAR Outlier DetectioN Using Diverse Synthetic Data

@@ Abstract
LiDAR-based semantic scene understanding is an important module in the modern autonomous driving perception stack. However, identifying outlier points in a LiDAR point cloud is challenging as LiDAR point clouds lack semantically-rich information. While former SOTA methods adopt heuristic architectures, we revisit this problem from the perspective of Selective Classification, which introduces a selective function into the standard closed-set classification setup. Our solution is built upon the basic idea of abstaining from choosing any inlier categories but learns a point-wise abstaining penalty with a margin-based loss. Apart from learning paradigms, synthesizing outliers to approximate unlimited real outliers is also critical, so we propose a strong synthesis pipeline that generates outliers originated from various factors: object categories, sampling patterns and sizes. We demonstrate that learning different abstaining penalties, apart from point-wise penalty, for different types of (synthesized) outliers can further improve the performance. We benchmark our method on SemanticKITTI and nuScenes and achieve SOTA results. Codes are available at https://github.com/Daniellli/LiON/.
@@ Introduction
LiDAR outlier detection (Li and Dong 2023) complements LiDAR semantic segmentation (Wang et al. 2024), aiming to enhance the model’s ability to recognize outliers without compromising its inlier segmentation performance.
This task is important and practical. Traditional segmentation methods (Wang et al. 2024; Li, Shum, and Breckon 2024) assume that the samples in the training and test sets belong to the same set of categories. Thus, these models are trained on inlier categories and tend to classify all inputs into one of the inlier categories. However, this assumption fails in real-world scenarios where outliers are present. For example, as shown in Fig. 1-(a), the model may randomly classify furniture that has not been seen in training, leading to disastrous consequences in the downstream planning stage.
Figure 1: (a). The semantic segmentation model fails to identify furniture because the training set does not include such objects; (b). Comparison of our ShapeNet outlier synthesis method and the former resize outlier synthesis method.
While 2D outlier detection has made significant strides, including optimizing inlier prediction (Tian et al. 2022; Liu et al. 2023; Miao et al. 2024), addressing outlier class imbalance (Choi, Jeong, and Choi 2023), evolving from pixel-wise to mask-based outlier detection methods (Zhang et al. 2024; Rai et al. 2023; Nayal et al. 2023; Zhang et al. 2024), developing promptable outlier detectors (Zhao et al. 2024; Li et al. 2024; Zhou et al. 2022), and utilizing model ensembling (Liu et al. 2024), the field of LiDAR outlier detection is still in its early stages (Li and Dong 2023). Seminal work REAL (Cen et al. 2022) proposes randomly choosing and resizing objects existing in the scene to synthesize outliers to approximate unlimited real outliers. For example, as shown in the left panel of Fig. 1-(b), the car is chosen and shrunk. This is viable but fail to represent the long-tail distribution of real outliers, in two regards. Firstly, objects from existing road scene understanding datasets are limited in category. Secondly, naive resizing violates the sampling pattern of real LiDAR sensors: points on enlarged objects get sparser while those on shrunk objects get denser. This leads to a shortcut problem: the model may find a trivial solution to distinguish outliers from inliers solely using point sparsity.
Besides, REAL empirically finds 2D outlier detection methods (Hendrycks and Gimpel 2018; Hendrycks et al. 2019; Gal and Ghahramani 2016) perform poorly in the 3D domain, as a large number of outliers are predicted as inliers with high confidence scores . To alleviate this phenomenon, in addition to the Cross Entropy (CE) loss, REAL designs a Calibration Cross Entropy(CCE) loss to calibrate the outlier probabilities in inlier prediction. However, we empirically find that CCE alleviates this at the cost of accuracy in predicting inliers. As shown in the right panel of Fig. 4, with CCE, the outlier probabilities of inlier points, which account for the vast majority of the total points, become very high (higher than 0.1), which is undesirable.
To address these issues, in this work, we propose a novel method LiON, aiming to mitigate the lack of semantically-rich information in LiDAR point clouds for outlier detection. We contribute from two perspectives: learning and data.
Learning. We reformulate the LiDAR outlier detection problem by applying Selective Classification (SC) principles (Feng et al. 2023; Chow 1970) and introduce a point-wise abstaining penalty learning paradigm to address the problem of unclear distinction in point clouds. While inspired by SC, our method differs significantly from SC in our point-wise design. Specifically, a diverse calibration factor is learned in a point-wise manner to more effectively capture subtle differences within a point cloud and calibrate the relationship between inlier and outlier classifiers. As a result, we mitigate the unclear distinction in point clouds caused by their lack of semantically-rich information by learning a diverse factor in a point-wise manner.
Data. Inspired by outlier exposure (Hendrycks, Mazeika, and Dietterich 2019), we propose introducing objects from an external dataset, ShapeNet (Chang et al. 2015), into existing scenes to synthesize outliers. ShapeNet, with its wide spectrum of categories and diverse geometries, can compensate for the long-tail distribution of real outliers. To ensure the realism of synthesized outliers, we take the LiDAR sampling pattern into consideration when merging randomly selected ShapeNet objects into road scenes. As illustrated in the right of Fig. 1-(b), our synthesized outliers are precisely aligned with objects in the scene with respect to point sparsity and occlusion. In this way, we mitigate the lack of semantically-rich information in LiDAR point clouds by utilizing realistic, ShapeNet outliers with diverse geometries.
[…]
@@
[…]
  • We utilize ShapeNet objects to synthesize realistic and diverse outliers to approximate unlimited real outliers.
@@
  • Our method has achieved new SOTA performance for LiDAR outlier detection not only in previously established outlier detection metrics, but also in the risk-coverage curve metric, on SemanticKITTI and NuScenes.
Related Works
Outlier Detection in Autonomous Driving. Outlier detection is vital for ensuring the safety of ego-cars in open-world environments by identifying outlier objects. Extensive research has been conducted in 2D perception, specifically with semantic segmentation models. Unsupervised methods (Jung et al. 2021; Bevandić et al. 2021) involve post-processing predicted logits from frozen segmentation models to detect outliers. Supervised methods (Tian et al. 2022; Grcić, Bevandić, and Šegvić 2022; Chan, Rottmann, and Gottschalk 2021) utilize auxiliary datasets like COCO (Lin et al. 2014) to synthesize outlier objects in training images (e.g., Cityscapes (Cordts et al. 2016)) and retrain the segmentation model using outlier exposure (Hendrycks, Mazeika, and Dietterich 2019; Zhou et al. 2024).
While significant advancements have been made in 2D outlier detection, the exploration in the context of LiDAR point clouds remains limited. Cen et al. (2022) use resize synthesis pipeline and calibration loss to achieve the discrimination of outlier points. However, their focus primarily revolves around the open-world segmentation setting without extensively analyzing calibration effectiveness. Li and Dong (2023) propose an adversarial prototype framework that improves performance but involves complex network design and computationally expensive training. Considering these limitations, we present our method and substantiate its effectiveness through extensive experiments.
Selective Classification. SC can be broadly classified into two groups: 1) the first group focuses on addressing SC through the use of additional heads/logits (Geifman and El-Yaniv 2017, 2019; Liu et al. 2019; Feng et al. 2023; Gal and Ghahramani 2016; Chow 1970); 2) the second group tackles SC through cost-sensitive classification techniques (Charoenphakdee et al. 2021; Mozannar and Sontag 2020). Motivated by the extra head/logits design, resembling outlier detection, we design a new outlier detector from the perspective of SC and utilize SC’s evaluation metrics as alternative performance measures for our outlier detector.
Synthetic Data. Due to the limited availability and high cost associated with acquiring high-quality real-world datasets, synthetic data has been extensively employed in the field of scene understanding (Su et al. 2015; Movshovitz-Attias, Kanade, and Sheikh 2016; Zhang et al. 2017b; Handa et al. 2016; McCormac et al. 2017; Zhang et al. 2017a; Song et al. 2017; Gao et al. 2021, 2024a, 2024b, 2023; Xu et al. 2024; Ding et al. 2024). In this work, we introduce ShapeNet objects into the existing scene to effectively compensate for the lack of semantically-rich information in point clouds.
Figure 2: Method pipeline: a point cloud containing outliers synthesized by ShapeNet is processed by a feature extractor to obtain features. These features are then used by inlier and outlier classifiers to predict class logits.
Preliminaries: Selective Classification
@@
We first formalize the definition of SC and put different methods under a unified lens. Experimentally, the risk-coverage trade-off analysis that comes along with this framework, allows us to reveal in-depth differences between methods.
@@
Definition. In SC, our goal is to learn predictive models that know what they do not know or when they should abstain [^1] The abstaining penalty will be defined later. from making decisions. Here we consider a generic SC definition, which is agnostic of network and application. A supervised classification task is formulated as follows. Let $\mathcal{X}$ be any feature space and $\mathcal{Y}$ a label space. In LiDAR outlier detection, $\mathcal{X}$ could be point clouds, and $\mathcal{Y}$ could be class labels [^2] These class labels include an outlier label. of each point cloud. Let $P(\mathcal{X},\mathcal{Y})$ be a distribution over $\mathcal{X}\times\mathcal{Y}$. A model $f:\mathcal{X}\rightarrow\mathcal{Y}$ is called a prediction function and its true risk used to evaluate the performance of $f$ w.r.t. $P$ is $R(f):=E_{P(\mathcal{X},\mathcal{Y})}[\ell(f(\mathbf{x}),\mathbf{y})]$, where $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a given loss function, for example, the Cross Entropy (CE) loss. Given a labeled set $S_{m}=\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{m}$ sampled i.i.d. from $P(\mathcal{X},\mathcal{Y})$, the empirical risk of the classifier $f$ is $\hat{r}(f|S_{m}):=\frac{1}{m}\sum_{i=1}^{m}\ell(f(\mathbf{x_{i}}),\mathbf{y_{i}})$.
@@
Finally, we clarify that the continuous risk and coverage defined above are calculated using a fixed set in practice. For any given labeled set $S_{m}$, the empirical selective risk is
@@
Remark. Later we will put methods under the unified lens of SC and use the risk-coverage curve to analyze them.
@@
LiDAR semantic segmentation is the task of assigning a class label from a predefined class label set to each point in a given point cloud. LiDAR outlier detection, on the other hand, is an extension to semantic segmentation, which aims to identify points that do not belong to the predefined inlier label set. While the seminal work proposes a viable solution REAL, we revisit this issue through the unified lens of SC and propose a solution that effectively mitigates the problem of unclear distinctions in point clouds. Apart from the learning paradigm, we design a novel outlier synthesis pipeline that leverages the richness of the ShapeNet and adheres to the realistic LiDAR distribution to compensate for the lack of semantically-rich information in point clouds.
Network Architecture Overview
As shown in Fig. 2, the input point cloud $\mathbf{x}\in\mathbb{R}^{n\times 3}$, sampled from $S_{m}$, is denoted on the left with ShapeNet objects integrated into it. Then, $\mathbf{x}$ is fed into the feature extractor $\psi$ followed by an inlier classifier $f$ to predict the inlier logit $\mathbf{\hat{y}}\in\mathbb{R}^{n\times c}$, where $c$ is the number of inlier classes. An outlier classifier $g$ is used to predict the outlier logit $\mathbf{\hat{o}}\in\mathbb{R}^{n\times 1}$. As such, $(f,g)$ instantiates a selective model mentioned above. These operations can be expressed as follows:
$\begin{gathered}\mathbf{\hat{y}}:=f(\psi(\mathbf{x}))\quad\mathbf{\hat{o}}:=g(\psi(\mathbf{x}))\\ \mathbf{\tilde{y}}:=[\mathbf{\hat{y}},\mathbf{\hat{o}}]:=\Biggl{\{}\mathbf{\tilde{y}_{i}}:=[\hat{y}_{i},\hat{o}_{i}]\Big{|}i=1,\ldots,n\Biggr{\}}\\ \mathbf{p}:=\Biggl{\{}p_{i,j}=\frac{e^{\tilde{y}_{i,j}}}{\sum_{k=1}^{c+1}e^{\tilde{y}_{i,k}}}\Big{|}i=1,\ldots,n;j=1,\ldots,c+1\Biggr{\}}\\ \mathbf{p^{y}},\mathbf{p^{o}}:=\mathbf{p}\end{gathered}$ (6)
@@
Cen et al. (2022) introduces the first LiDAR outlier detector, REAL. Their added dummy classifiers can also be thought of as $g$, under the SC framework, but there is a key difference. They observe that numerous real outliers are wrongly classified as inlier classes with high probabilities. To address this issue, they propose a Calibration Cross Entropy (CCE) loss function to drive the outlier logit of the inlier sample to the second largest. We formalize [^3] This is equivalent to the loss used in the REAL paper although REAL describes it in a different manner. this loss as follows:
$\begin{gathered}\ell:=\frac{1}{m}\sum_{S_{m}}\frac{1}{n}\sum_{i=1}^{n}\Biggl{\{}\underbrace{-\log{p_{i,y_{i}}}}_{\text{CE}}\\ \underbrace{-\lambda\mathbb{I}(y_{i}\neq c+1)\log{\frac{e^{\tilde{y}_{i,c+1}}}{\sum_{k=1\&k\neq y_{i}}^{c+1}e^{\tilde{y}_{i,k}}}}}_{\text{CCE}}\Biggr{\}}\\ \end{gathered}$ (7)
Here, $y_{i}\in\{1,\ldots,c,c+1\}$ signifies the ground truth corresponding to $x_{i}\in\mathbf{x}$, where $\{1,\ldots,c\}$ represents inlier class labels while $\{c+1\}$ indicates the outlier class label. $\mathbf{y}$ and $\mathbf{x}$ are sampled i.i.d. from $S_{m}$. $\mathbb{I}(\cdot)$ is the indicator function. $\lambda$ is a hyperparameter. A notable fact is that REAL does not provide an ablation study for this CCE loss and as shown in  Tab. 1, removing this CCE loss can indeed improve standard outlier detection metrics. But why this happens cannot be understood through metrics like AUPR or AUROC, highlighting the need to revisit REAL under the lens of SC.
@@
CE CCE AUPR AUROC $\text{mIoU}_{\text{old}}$ 20.00 84.90 57.80 26.68 87.60 58.28
Table 1: Removing CCE in REAL can actually improve the outlier detection metrics on SemanticKITTI.
Figure 3: Our outlier synthesis pipeline: (a) loading a ShapeNet object; (b) randomly moving it away from the scene center; (c) randomly rotating it around the gravity direction; (d) randomly resizing it; (e) putting it on ground; (f) resampling points on the object to blend into the scene. We repeat this pipeline on the fly $G$ times for inserting $G$ objects.
Point-wise Abstaining Penalty Learning
Figure 4: Statistics of outlier probability $\mathbf{p^{o}}$ for inlier and outlier samples under different settings of REAL on SemanticKITTI. For inliers, we would like to observe more samples with $\mathbf{p^{o}}\in[0,0.1]$, while for outliers, less samples with $\mathbf{p^{o}}\in[0,0.1]$ is desirable.
Point-wise upgrade. The reason why the disadvantages of CCE exceed its benefits is its sub-optimal calibration for the inlier probabilities $\mathbf{p^{y}}$ and the outlier probabilities $\mathbf{p^{o}}$. Inspired by ‘learning to abstain’ (Liu et al. 2019), where the calibration between the reject and non-reject options is adeptly handled, we propose our new learning paradigm. Specifically, Liu et al. (2019) employs a fixed calibrating factor to achieve this. While this may suffice for image-level SC, it is inadequate for point-wise outlier detection. This is because, in LiDAR outlier detection, we need a different calibrating factor for each point to capture the subtle differences between them, such as between remote (sparse) and near (dense) points, as well as between inlier and outlier points. Therefore, we upgrade this fixed calibrating factor to a point-wise one, defined as abstaining penalty $\alpha$, and introduce a point-wise penalty loss to supervise the network to learn the subtle differences between various points, which can be expressed as follows:
$\begin{gathered}\mathbf{\alpha}:=\left\{\alpha_{i}=-\log(\sum_{j=1}^{c}e^{\hat{y}_{i,j}})|i=1,\ldots,n\right\}\\ \ell^{\text{penalty}}:=\frac{1}{m}\sum_{S_{m}}\frac{1}{n}\sum_{i=1}^{n}\Biggl{\{}\mathbb{I}(y_{i}\neq c+1)\text{max}(\alpha_{i}-m_{\text{in}},0)\\ +\mathbb{I}(y_{i}=c+1)\text{max}(m_{\text{out}}-\alpha_{i},0)\Biggr{\}}\end{gathered}$ (8)
As shown in Fig. 5-(c), the hyperparameters $m_{\text{in}}$ and $m_{\text{out}}$ ensure that the inliers are associated with penalties lower than $m_{\text{in}}$, while the outliers exhibit penalties higher than $m_{\text{out}}$. Note that in our experiments, the penalties are negative and we set the value of $m_{\text{in}}$ to -12, and $m_{\text{out}}$ to -6.
Moreover, our new ‘learning to abstain’ formulation for this task with this point-wise penalty is defined as follows:
@@
Intuition. Minimizing the point-wise penalty loss Eq. 8 results in assigning lower $\alpha_{i}$ to inliers, thereby leading to higher values of $(-\alpha_{i})^{2}$. This effectively suppresses the contribution of $p^{o}_{i}$ and allows $p^{y}_{i,y_{i}}$ to play a dominant role in the point-wise abstain loss Eq. 9. Likewise, higher $\alpha_{i}$ are allocated to outliers, resulting in lower values of $(-\alpha_{i})^{2}$. This allows $p^{o}_{i}$ to play a dominant role and, consequently, suppresses the contribution of $\{p^{y}_{i,j}|j=1,\ldots,c\}$ in the point-wise abstain loss Eq. 9. Since this learning paradigm is defined in a point-wise manner, it has the potential to capture the subtle difference between inliers and outliers despite LiDAR point clouds lack semantically-rich information.
The total loss can be expressed as follows: